40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Mon Nov 9 15:49:58 2020, Aidan, Update, 2um Photodiodes, JPL PD resurrection (cont.) 
    Reply  Tue Nov 10 12:46:34 2020, Aidan, Update, 2um Photodiodes, JPL PD resurrection (cont.) temperature_diff_shield_v_PD.png
Message ID: 2516     Entry time: Mon Nov 9 15:49:58 2020     Reply to this: 2517
Author: Aidan 
Type: Update 
Category: 2um Photodiodes 
Subject: JPL PD resurrection (cont.) 

Okay - all the steps in the procedure of eLOG 2476 have been verified as working - with the exception of the RTDs in the chamber.

 With regards to taking dark noise spectra at different biases and temperatures, looks like Raymond took spectra with biases of [50, 100, 200, 400, 600, 1000]mV. If no objections, I’ll stick to that number of measurements. 

I’m a bit pushed for time with other stuff. I wonder if the shield RTD is sufficient to run tests on the system? I’ll go back through the data and see how reproducible the relationship between shield temperature and PD temperature is. If it is reliable then in the interests of time, I’m going to forgo re-installing the extra RTDs in the chamber just now.

 

ELOG V3.1.3-